### v2

# The universal family of punctured Riemann surfaces is Stein

Abstract.
We show that for any pair of integers $n\geq 1$ and $g\geq 0$ the universal family of $n$-punctured compact Riemann surfaces of genus $g$ is a Stein manifold.
Key words and phrases:
Riemann surface, Teichmüller space, universal family, Stein manifold, Oka manifold
2020 Mathematics Subject Classification:
Primary 32G15; secondary 32Q28, 32Q56
1. Introduction
[…]
(1.1) $V(g,n)=\widehat{V}(g,n)\setminus\bigcup_{i=1}^{n}s_{i}(T(g,n))$
of $\widehat{V}(g,n)$ is the universal family of $n$-punctured compact Riemann surfaces of genus $g$.
In this paper we prove the following result.
Theorem 1.1.
If $n\geq 1$ then the universal family $V(g,n)$ is a Stein manifold.
The theorem is trivial if $g=0$ and $n\in\{1,2,3\}$ since $T(0,n)$ is then a singleton and $V(0,n)$ equals $\mathbb{C}$, $\mathbb{C}^{*}=\mathbb{C}\setminus\{0\}$, and $\mathbb{C}\setminus\{0,1\}$, respectively. If $2g+n\geq 3$ then $T(g,n)$ is biholomorphic to a bounded topologically contractible Stein domain in $\mathbb{C}^{3g-3+n}$ [22, p. 161], but it cannot be holomorphically realised as a convex domain in $\mathbb{C}^{3g-3+n}$ if $g\geq 2$ (see Marković [21]). Note also that $V(g,0)$ is holomorphically convex and all holomorphic functions on it come from the base $T(g,0)$.
The $n$ holomorphic sections $s_{1},\ldots,s_{n}$ of the submersion $\pi:\widehat{V}(g,n)\to T(g,n)$ in (1.1), corresponding to the punctures of $V(g,n)$, are called canonical sections in the literrature. The problem of describing the global holomorphic sections of this fibration was first considered by Hubbard, who showed in [19] that $V(g,0)\to T(g,0)$ has no holomorphic sections if $g\geq 3$ and six sections if $g=2$. See also [20]. Earle and Kra proved in [7, p. 50] that for any $n\geq 1$, $\pi:\widehat{V}(g,n)\to T(g,n)$ has exactly $n$ sections if $g\geq 3$ and exactly $2n+6$ sections if $g=2$.
Theorem 1.1 has interesting consequences for function theory and complex geometry. By classical results of Remmert, Bishop, and Narasiman (see [9, Theorem 2.4.1] and the references therein), the Stein manifold $V(g,n)$ for $n\geq 1$ admits a proper holomorphic embedding in $\mathbb{C}^{N}$ with $N=2\dim V(g,n)+1$ (which equals 6g6+2n+16g-6+2n+1 if $2g+n\geq 3$). When $\dim V(g,n)\geq 2$, it also embeds properly holomorphically in $\mathbb{C}^{N}$ with $N=\big[\frac{3\dim V(g,n)}{2}\big]+1$ by the optimal embedding theorems of Eliashberg and Gromov [8] and Schürmann [24]. Since every fibre of $V(g,n)$ is a noncompact Riemann surface of finite conformal type, and hence is biholomorphic to a closed affine algebraic curve in $\mathbb{C}^{3}$, the following question is natural.
Problem 1.2.
Does $V(g,n)$ for $n\geq 1$ admit a proper holomorphic embedding in some $\mathbb{C}^{N}$ which is algebraic on every fibre?
Except in the trivial cases, $T(g,n)$ is not affine algebraic, and it is not reasonable to expect that $V(g,n)$ is affine algebraic.
Since $T(g,n)$ is contractible, the holomorphic submersion $\pi:V(g,n)\to T(g,n)$ is smoothly trivial and the inclusion of any fibre of $\pi$ in $V(g,n)$ is a homotopy equivalence. It follows that for any manifold $Y$, the restriction of a continuous map $V(g,n)\to Y$ to any fibre of $V(g,n)$ lies in the same homotopy class of maps $M\to Y$, where $M$ is the underlying $n$-punctured smooth compact surface of genus $g$. The following is an immediate corollary to this observation, Theorem 1.1, and the main result of Oka theory [9, Theorem 5.4.4]. For Oka manifolds, see [9, Chap. 5] and the surveys [10, 12, 13].
Corollary 1.3.
Let $\pi:V(g,n)\to T(g,n)$ be as above, $n\geq 1$, and let $Y$ be an Oka manifold. There is a holomorphic map $V(g,n)\to Y$ in every homotopy class. Furthermore, a holomorphic map $M_{t}\to Y$, $t\in T(g,n)$, from any fibre of $V(g,n)$ extends to a holomorphic map $V(g,n)\to Y$. More generally, given a closed complex subvariety $T^{\prime}$ of $T(g,n)$, every continuous map $F_{0}:V(g,n)\to Y$ which is holomorphic on $\pi^{-1}(T^{\prime})$ is homotopic to a holomorphic map $F:V(g,n)\to Y$ by a homotopy which is fixed on $\pi^{-1}(T^{\prime})$.
If $n\geq 1$ and $(g,n)\neq(0,1)$ then $V(g,n)$ is not simply connected, and its homotopy type is that of a finite bouquet of circles. In this case, homotopically nontrivial maps $V(g,n)\to Y$ exist whenever the manifold $Y$ is not simply connected. This gives the following corollary. The last statement follows by taking the Oka manifold $Y=\mathbb{C}^{*}$.
Corollary 1.4.
If $n\geq 1$, $(g,n)\neq(0,1)$, and $Y$ is an Oka manifold which is not simply connected, there is a holomorphic map $V(g,n)\to Y$ which is nonconstant on every fibre. In particular, $V(g,n)$ admits a nowhere vanishing holomorphic function which is nonconstant on every fibre.
Another way to obtain fibrewise nonconstant holomorphic maps from $V(g,n)$ to an Oka manifold is to inductively use the Oka property with approximation on compact holomorphically convex subsets of the Stein manifold $V(g,n)$; see [9, Theorem 5.4.4]. The Oka principle can be used to obtain many further properties of holomorphic universal families $V(g,n)\to Y$ in any Oka manifold.
In this connection, we mention that in [11] the Oka principle was established for families of maps from very general families of open Riemann surfaces $(M,J_{t})$ to any Oka manifold $Y$, where $M$ is a smooth open surface and $J_{t}$ are complex structures of some local Hölder class on $M$ depending continuously or smoothly on the parameter $t\in T$ in a suitable topological space, which is not a complex manifold in general. The Riemann surfaces in such families need not belong to the same Teichmüller space. For example, a punctured Riemann surface can be a member of a family in which the punctures develop into boundary curves and vice versa. The $J_{t}$-holomorphic maps $F_{t}:(M,J_{t})\to Y$ $(t\in T)$ furnished by [11, Theorem 1.6] are merely continuous or smooth in the parameter. In [14] these results were extended to maps from tame families of Stein manifolds of arbitrary dimension to Oka manifolds. Here, a continuous family $\{J_{t}\}_{t\in T}$ of Stein structures on a smooth manifold $X$ said to be tame if the holomorphically convex hulls of any compact subset $K\subset X$ are upper semicontinuous with respect to the parameter $t\in T$. This is always the case in a family of open Riemann surfaces but it fails general in higher dimensions, even on affine spaces.
2. Proof of Theorem 1.1
In view of the description of the Teichmüller submersion $\pi:V(g,n)\to T(g,n)$ in (1.1), Theorem 1.1 is an immediate consequence of the following more general result with arbitrary Stein manifold as the base.
Theorem 2.1.
Assume that $X$ is a Stein manifold, $Z$ is a complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion, and $s_{1},\ldots,s_{n}:X\to Z$ are holomorphic sections with pairwise disjoint images for some $n\geq 1$. Then, the domain $\Omega=Z\setminus\bigcup_{i=1}^{n}s_{i}(X)$ is Stein.
The conclusion fails if the fibres of $\pi$ have complex dimension >1>1 or the sections $s_{i}$ are not holomorphic. In such case, the domain $\Omega$ in the theorem fails to be locally pseudoconvex at some boundary point $s_{i}(x)$, $x\in X$. Note that Stein complements of smooth complex hypersurfaces in compact Kähler manifolds have recently been studied by Höring and Peternell [18] where the reader can find references to earlier works. In our case, $Z$ is not compact unless $X$ is a point.
Proof.
Let $\pi:Z\to X$ be as in the theorem. We may assume that $X$ is connected. Note that every fibre $Z_{x}=\pi^{-1}(x)$, $x\in X$, is a compact Riemann surface, and the fibres are diffeomorphic but not necessarily biholomorphic to each other. Hence, $\{Z_{x}\}_{x\in X}$ is a holomorphic family of compact Riemann surfaces and $\Omega_{x}=Z_{x}\setminus\bigcup_{i=1}^{n}s_{i}(x)$ $(x\in X)$ is a holomorphic family of $n$-punctured Riemann surfaces. Each $H_{i}=s_{i}(X)$ is a closed complex hypersurface in $Z$ whose ideal sheaf is a principal, that is, it is locally near each point of $H_{i}$ generated by a single holomorphic function.
Recall the following result (see Grauert and Remmert [16, Theorem 5, p. 129]): If $Z$ is a Stein space and $H$ is a closed complex analytic hypersurface in $Z$ (of pure codimension one) whose ideal sheaf is a principal ideal sheaf, then $Z\setminus H$ is also Stein. If $Z$ is nonsingular (a complex manifold) then the ideal sheaf of any closed complex hypersurface in $Z$ is a principal ideal sheaf (see [16, Chap. A.3.5]). Hence, is suffices to prove the theorem in the case $n=1$, that is, to show that the complement $Z\setminus s(X)$ of a single holomorphic section $s:X\to Z$ is a Stein manifold.
By a theorem of Siu [25], the Stein hypersurface $H=s(X)$ has a basis of open Stein neighbourhoods $U$ in $Z$. Since $U\setminus H$ is a Stein manifold by the aforementioned theorem [16, Theorem 5, p. 129], it admits a strongly plurisubharmonic exhaustion function $\phi:U\setminus H\to\mathbb{R}_{+}$. To prove the theorem, we shall construct a strongly plurisubharmonic exhaustion function $Z\setminus H\to\mathbb{R}_{+}$; a theorem of Grauert [15] will then imply that $Z\setminus H$ is Stein.
Fix a point $x_{0}\in X$ and set $z_{0}=s(x_{0})\in H\subset Z$. Since $\pi:Z\to X$ is a holomorphic submersion with compact one-dimensional fibres, it is a smooth fibre bundle whose fibre $M$ is a compact smooth surface. In particular, there is a neighbourhood $X_{0}\subset X$ of $x_{0}$ such that the restricted bundle $Z|X_{0}=\pi^{-1}(X_{0})\to X_{0}$ can be smoothly identified with the trivial bundle $X_{0}\times M\to X_{0}$. In this identification, $z_{0}=(x_{0},p_{0})$ with $p_{0}\in M$. Since $\phi$ tends to $\infty$ along $H$, there are small smoothly bounded open discs $D\Subset D^{\prime}\Subset M$ with $p_{0}\in D$ such that
(2.1) $\inf_{p\in bD}\phi(x_{0},p)>\max_{p\in bD^{\prime}}\phi(x_{0},p).$
The set $O=M\setminus D$ is a compact bordered Riemann surface with smooth boundary $bO=bD$, endowed with the complex structure inherited by the identification $M\cong Z_{x_{0}}=\pi^{-1}(x_{0})$. Note that $bD^{\prime}$ is contained in the interior of $O$. It follows from (2.1) and standard results that there is a smooth strongly subharmonic function $u_{0}:O\to\mathbb{R}_{+}$ such that
$u_{0}<\phi(x_{0},\cdotp)$ on $bO=bD$ and $u_{0}>\phi(x_{0},\cdotp)$ on $bD^{\prime}$.
Shrinking the neighbourhood $X_{0}\subset X$ of $x_{0}$ if necessary, the following conditions hold for every $x\in X_{0}$, where we use the smooth fibre bundle isomorphism $Z|X_{0}\cong X_{0}\times M$:
  1. (a)
    $s(x)\in D$,
  2. (b)
    the function $u(x,\cdotp)=u_{0}$ is strongly subharmonic on $O$ in the complex structure on $Z_{x}\cong M$,
  3. (c)
    $u(x,\cdotp)<\phi(x,\cdotp)$ on $bO=bD$, and
  4. (d)
    $u(x,\cdotp)>\phi(x,\cdotp)$ on $bD^{\prime}$.
Condition (b) holds since being strongly subharmonic on a compact subset is a stable property under small smooth deformations of the complex structure. We define a function $\rho_{0}:(X_{0}\times M)\setminus H\to\mathbb{R}_{+}$ by taking for every $x\in X_{0}$:
$\rho_{0}(x,p)=\begin{cases}\phi(x,p),&p\in D\setminus\{s(x)\};\\ \max\{\phi(x,p),u(x,p)\},&p\in D^{\prime}\setminus D;\\ u(x,p),&p\in M\setminus D^{\prime}.\end{cases}$
Note that $\rho_{0}$ is well defined, piecewise smooth, strongly subharmonic on each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X_{0})$, and it agrees with $\phi$ on $(X_{0}\times D)\setminus H$. By using the regularised maximum in the definition of $\rho_{0}$ (see [9, Eq. (3.1), p. 69]), we may assume that $\rho_{0}$ is smooth and enjoys the other stated properties.
This construction gives an open locally finite cover $\{X_{j}\}_{j=1}^{\infty}$ of $X$ with smooth fibre bundle trivialisations $Z|X_{j}\cong X_{j}\times M$, discs $D_{j}\subset M$ such that $s(x)\in D_{j}$ for all $x\in X_{j}$, and smooth functions $\rho_{j}:(Z|X_{j})\setminus H=\pi^{-1}(X_{j})\setminus H\to\mathbb{R}_{+}$ such that $\rho_{j}$ is strongly subharmonic on each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X_{j})$ and it agrees with $\phi$ on $(X_{j}\times D_{j})\setminus H$. Let $\{\chi_{j}\}_{j}$ be a smooth partition of unity on $X$ with compact supports $\mathrm{supp}(\chi_{j})\subset X_{j}$ for each $j$. Set
$\rho=\sum_{j=1}^{\infty}\chi_{j}\rho_{j}:Z\setminus H\to\mathbb{R}_{+}.$
By the construction, the restriction of $\rho$ to each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X)$ is strongly subharmonic, and there is an open neighbourhood $U_{0}\subset U\subset Z$ of $H$ such that $\rho=\phi$ holds on $U_{0}\setminus H$. In particular, $\rho$ is strongly plurisubharmonic on $U_{0}\setminus H$.
Note that for every compact set $K\subset X$, the set $\pi^{-1}(K)\setminus U_{0}\subset Z\setminus H$ is compact. Hence, by choosing a strongly plurisubharmonic exhaustion function $\tau:X\to\mathbb{R}_{+}$ whose Levi form $dd^{c}\tau$ grows fast enough, we can ensure that $\rho+\tau\circ\pi:Z\setminus H\to\mathbb{R}_{+}$ is a strongly plurisubharmonic exhaustion function. Indeed, denoting by $J$ the almost complex structure operator on $Z$ and by $d^{c}$ the conjugate differential defined by $(d^{c}\rho)(z,\xi)=-d\rho(z,J\xi)$ for $z\in Z$ and $\xi\in T_{z}Z$, $\rho$ is strongly plurisubharmonic at $z\in Z$ if and only if $(dd^{c}\rho)(z,\xi\wedge J\xi)>0$ for every tangent vector $0\neq\xi\in T_{z}Z$. (Up to a positive factor this equals the Laplace of $\rho$ on the 2-plane $\mathrm{span}(\xi,J\xi)\subset T_{z}Z$.) Since the function $\rho$ constructed above is strongly subharmonic on every fibre $Z_{x}\setminus\{s(x)\}$, $x\in X$, we have that
$(dd^{c}\rho)(z,\xi\wedge J\xi)>0\ \ \text{if $z\in Z\setminus H$ and $0\neq\xi\in\ker d\pi_{z}$.}$
Hence, the eigenvectors of the Levi form $(dd^{c}\rho)(z,\cdotp)$ associated to non-positive eigenvalues lie in a closed cone $C_{z}\subset T_{z}Z$ which intersects $\ker d\pi_{z}$ only in the origin. It follows that if $\tau:X\to\mathbb{R}$ is such that $dd^{c}\tau>0$ is sufficiently big on $T_{x}X$, $x=\pi(z)$, then $dd^{c}\rho+dd^{c}(\tau\circ\pi)>0$ on $T_{z}Z$. Furthermore, the estimates are uniform on the compact set $\pi^{-1}(K)\setminus U_{0}$, where $U_{0}\subset Z$ is a neighbourhood of $H$ as above such that $dd^{c}\rho>0$ on $U_{0}\setminus H$. To see that $\tau$ can be chosen such that $dd^{c}\tau$ grows as fast as desired, note that if $h:\mathbb{R}\to\mathbb{R}$ is a $\mathscr{C}^{2}$ function then for each point $x\in X$ and tangent vector $\xi\in T_{x}X$ we have that
$dd^{c}(h\circ\tau)(x,\xi\wedge J\xi)=h^{\prime}(\tau(x))\,(dd^{c}\tau)(x,\xi\wedge J\xi)+h^{\prime\prime}(\tau(x))\,\big(|d\tau(x,\xi)|^{2}+|d\tau(x,J\xi|^{2}\big).$
Hence, if $\tau$ is a strongly plurisubharmonic exhaustion function on $X$ and the function $h:\mathbb{R}\to\mathbb{R}$ is chosen such that $h^{\prime\prime}\geq 0$ and $h^{\prime}$ grows sufficiently fast, then $dd^{c}(h\circ u)$ also grows as fast as desired.
This completes the proof of Theorem 2.1, and it also proves Theorem 1.1. ∎
A minor modification of the proof of Theorem 2.1 gives the following more general result.
Theorem 2.2.
Assume that $X$ is a connected Stein manifold, $Z$ is a complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion, and $H$ is a closed complex subvariety of $Z$ of pure codimension one which does not contain any fibre of $\pi$. Then, the domain $\Omega=Z\setminus H$ is Stein.
Sketch of proof.
The restricted projection $\pi:H\to X$ is a proper holomorphic map. Since $H$ is of pure codimension one and does not contain any fibre of $\pi$, $\pi:H\to X$ is a finite holomorphic map (a branched holomorphic cover). The proper mapping theorem of Remmert [23] (see also Chirka [6, p. 29]) implies that $\pi(H)$ is a closed complex subvariety of $X$ of pure dimension $n$, hence $\pi(H)=X$ since $X$ is connected, and $H$ is Stein by [16, Theorem 1 (d), p. 125]. By Siu’s theorem [25], $H$ admits a basis of open Stein neighbourhoods $U\subset Z$. By the argument in the proof of Theorem 2.1, the ideal sheaf of $H$ is a principal ideal sheaf, so $U\setminus H$ is Stein for any Stein neighbourhood $U$ of $H$ by [16, Theorem 5, p. 129]. The proof can now be completed by a similar argument as in the proof of Theorem 2.1, and we leave further details to the reader. ∎
Acknowledgements. Research was supported by the European Union (ERC Advanced grant HPDR, 101053085) and grants P1-0291 and N1-0237 from ARIS, Republic of Slovenia. I wish to thank Finnur Lárusson for posing the question answered by Theorem 1.1 in a private communication.

### v3

@@
 Abstract.
-We show that for any pair of integers $n\geq 1$ and $g\geq 0$ the universal family of $n$-punctured compact Riemann surfaces of genus $g$ is a Stein manifold.
+We show that for any pair of integers $g\geq 0$ and $n\geq 1$ the universal family of $n$-punctured compact Riemann surfaces of genus $g$ is a Stein manifold.
+We describe its basic function theoretic properties and pose several challenging questions.
 Key words and phrases:
@@
 Primary 32G15; secondary 32Q28, 32Q56
-1. Introduction
 […]
-(1.1) $V(g,n)=\widehat{V}(g,n)\setminus\bigcup_{i=1}^{n}s_{i}(T(g,n))$
+(1) $V(g,n)=\widehat{V}(g,n)\setminus\bigcup_{i=1}^{n}s_{i}(T(g,n))$
 of $\widehat{V}(g,n)$ is the universal family of $n$-punctured compact Riemann surfaces of genus $g$.
+If $2g+n\geq 3$ then the Teichmüller family $\pi:V(g,n)\to T(g,n)$ is the universal object in the complex analytic category of topologically marked holomorphically varying families of $n$-punctured genus $g$ Riemann surfaces (see [28, Theorem 5.4.3]).
 In this paper we prove the following result.
-Theorem 1.1.
+Theorem 1.
 If $n\geq 1$ then the universal family $V(g,n)$ is a Stein manifold.
+This is a special case of Theorems 8 and 9 given in the sequel.
 The theorem is trivial if $g=0$ and $n\in\{1,2,3\}$ since $T(0,n)$ is then a singleton and $V(0,n)$ equals $\mathbb{C}$, $\mathbb{C}^{*}=\mathbb{C}\setminus\{0\}$, and $\mathbb{C}\setminus\{0,1\}$, respectively.
-If $2g+n\geq 3$ then $T(g,n)$ is biholomorphic to a bounded topologically contractible Stein domain in $\mathbb{C}^{3g-3+n}$ [22, p.
+If $2g+n\geq 3$ then $T(g,n)$ is biholomorphic to a bounded topologically contractible Stein domain in $\mathbb{C}^{3g-3+n}$ [28, p.
 161], but it cannot be holomorphically realised as a convex domain in $\mathbb{C}^{3g-3+n}$ if $g\geq 2$ (see Marković [27]).
-Note also that $V(g,0)$ is holomorphically convex and all holomorphic functions on it come from the base $T(g,0)$.
-The $n$ holomorphic sections $s_{1},\ldots,s_{n}$ of the submersion $\pi:\widehat{V}(g,n)\to T(g,n)$ in (1.1), corresponding to the punctures of $V(g,n)$, are called canonical sections in the literrature.
+The $n$ holomorphic sections $s_{1},\ldots,s_{n}$ of the submersion $\pi:\widehat{V}(g,n)\to T(g,n)$ in (1), corresponding to the punctures of $V(g,n)$, are called canonical sections.
 The problem of describing the global holomorphic sections of this fibration was first considered by Hubbard, who showed in [25] that $V(g,0)\to T(g,0)$ has no holomorphic sections if $g\geq 3$ and six sections if $g=2$.
 See also [26].
-Earle and Kra proved in [7, p.
-50] that for any $n\geq 1$, $\pi:\widehat{V}(g,n)\to T(g,n)$ has exactly $n$ sections if $g\geq 3$ and exactly $2n+6$ sections if $g=2$.
-Theorem 1.1 has interesting consequences for function theory and complex geometry.
-By classical results of Remmert, Bishop, and Narasiman (see [9, Theorem 2.4.1] and the references therein), the Stein manifold $V(g,n)$ for $n\geq 1$ admits a proper holomorphic embedding in $\mathbb{C}^{N}$ with $N=2\dim V(g,n)+1$ (which equals 6g6+2n+16g-6+2n+1 if $2g+n\geq 3$).
-When $\dim V(g,n)\geq 2$, it also embeds properly holomorphically in $\mathbb{C}^{N}$ with $N=\big[\frac{3\dim V(g,n)}{2}\big]+1$ by the optimal embedding theorems of Eliashberg and Gromov [8] and Schürmann [24].
-Since every fibre of $V(g,n)$ is a noncompact Riemann surface of finite conformal type, and hence is biholomorphic to a closed affine algebraic curve in $\mathbb{C}^{3}$, the following question is natural.
-Problem 1.2.
+Earle and Kra proved in [10, p.
+50] that for $n\geq 1$, $\pi:\widehat{V}(g,n)\to T(g,n)$ has exactly $n$ sections if $g\geq 3$ and exactly $2n+6$ sections if $g=2$.
+Theorem 1 has interesting consequences for function theory and complex geometry.
+By classical results of Remmert, Bishop, and Narasiman (see [15, Theorem 2.4.1] and the references therein), the Stein manifold $V(g,n)$ for $n\geq 1$ admits a proper holomorphic embedding in $\mathbb{C}^{N}$ with $N=2\dim V(g,n)+1$, which equals 6g6+2n+16g-6+2n+1 if $2g+n\geq 3$.
+When $\dim V(g,n)\geq 2$, it also embeds properly holomorphically in $\mathbb{C}^{N}$ with $N=\big[\frac{3\dim V(g,n)}{2}\big]+1$ (see Eliashberg and Gromov [11] and Schürmann [31], or the exposition in [15, Secs.
+9.3–9.4]).
+Since every fibre of $V(g,n)$ is biholomorphic to a closed affine algebraic curve in $\mathbb{C}^{3}$, the following question is natural.
+Problem 2.
 Does $V(g,n)$ for $n\geq 1$ admit a proper holomorphic embedding in some $\mathbb{C}^{N}$ which is algebraic on every fibre?
@@
 It follows that for any manifold $Y$, the restriction of a continuous map $V(g,n)\to Y$ to any fibre of $V(g,n)$ lies in the same homotopy class of maps $M\to Y$, where $M$ is the underlying $n$-punctured smooth compact surface of genus $g$.
-The following is an immediate corollary to this observation, Theorem 1.1, and the main result of Oka theory [9, Theorem 5.4.4].
-For Oka manifolds, see [9, Chap.
+The following is an immediate corollary to this observation, Theorem 1, and the main result of Oka theory [15, Theorem 5.4.4].
+For Oka manifolds, see [15, Chap.
 5] and the surveys [13, 12, 17].
-Corollary 1.3.
+Corollary 3.
 Let $\pi:V(g,n)\to T(g,n)$ be as above, $n\geq 1$, and let $Y$ be an Oka manifold.
@@
 The last statement follows by taking the Oka manifold $Y=\mathbb{C}^{*}$.
-Corollary 1.4.
+Corollary 4.
 If $n\geq 1$, $(g,n)\neq(0,1)$, and $Y$ is an Oka manifold which is not simply connected, there is a holomorphic map $V(g,n)\to Y$ which is nonconstant on every fibre.
 In particular, $V(g,n)$ admits a nowhere vanishing holomorphic function which is nonconstant on every fibre.
-Another way to obtain fibrewise nonconstant holomorphic maps from $V(g,n)$ to an Oka manifold is to inductively use the Oka property with approximation on compact holomorphically convex subsets of the Stein manifold $V(g,n)$; see [9, Theorem 5.4.4].
+Another way to obtain fibrewise nonconstant holomorphic maps from $V(g,n)$ to an Oka manifold is to inductively use the Oka property with approximation on compact holomorphically convex subsets of the Stein manifold $V(g,n)$; see [15, Theorem 5.4.4].
 The Oka principle can be used to obtain many further properties of holomorphic universal families $V(g,n)\to Y$ in any Oka manifold.
-In this connection, we mention that in [11] the Oka principle was established for families of maps from very general families of open Riemann surfaces $(M,J_{t})$ to any Oka manifold $Y$, where $M$ is a smooth open surface and $J_{t}$ are complex structures of some local Hölder class on $M$ depending continuously or smoothly on the parameter $t\in T$ in a suitable topological space, which is not a complex manifold in general.
+The fact that $V(g,n)$ for $n\geq 1$ is homotopy equivalent to a bouquet of circles implies that every complex vector bundle on $V(g,n)$ is topologically trivial.
+Since $V(g,n)$ is Stein, the Oka–Grauert principle ([19] or [15, Theorem 3.2.1]) implies the following.
+Proposition 5.
+Every holomorphic vector bundle on $V(g,n)$ for $n\geq 1$ is holomorphically trivial.
+Corollary 6.
+Assume that $g\geq 0$ and $n\geq 1$.
+  1. (a)
+    There exists a nowhere vanishing holomorphic vector field $\xi$ on $V(g,n)$ which is tangent to the fibres of the projection $\pi:V(g,n)\to T(g,n)$, that is, $d\pi(\xi)=0$.
+  2. (b)
+    With $\xi$ as in (a), there exists a holomorphic $1$-form $\theta$ on $V(g,n)$ satisfying $\langle\theta,\xi\rangle=1$.
+In particular, $\theta$ is nowhere vanishing on the tangent bundle to any fibre of $\pi$.
+Proof.
+Part (a) follows by applying Proposition 5 to the holomorphic line bundle $\ker d\pi\to V(g,n)$, the vertical tangent bundle of the holomorphic submersion $\pi:V(g,n)\to T(g,n)$.
+To see (b), consider the following short exact sequence of vector bundles over $V(g,n)$:
+$0\longrightarrow\ker d\pi\lhook\joinrel\relbar\joinrel\rightarrow TV(g,n)\stackrel{{\scriptstyle\alpha}}{{\longrightarrow}}H:=TV(g,n)/\ker\pi\longrightarrow 0.$
+Here, $TV(g,n)$ denote the tangent bundle of $V(g,n)$.
+By Cartan’s Theorem B the sequence splits, i.e.
+there is a holomorphic vector bundle injection $\sigma:H\hookrightarrow TV(g,n)$ such that $\alpha\circ\sigma=\mathrm{Id}_{H}$.
+Hence, $TV(g,n)=\ker d\pi\oplus\sigma(H)=\mathbb{C}\xi\oplus\sigma(H)$, where $\xi$ is as in part (a).
+The unique holomorphic $1$-form $\theta$ on $V(g,n)$ satisfying $\langle\theta,\xi\rangle=1$ and $\xi=0$ on $\sigma(H)$ clearly satisfies part (b).
+∎
+By the Gunning–Narasimhan theorem [23], every open Riemann surface $M$ admits a holomorphic immersion $f:M\to\mathbb{C}$. In view of Corollary 6 (b), the following is a natural question.
+Problem 7.
+Let $n\geq 1$. Does there exist a holomorphic function $f:V(g,n)\to\mathbb{C}$ whose restriction to any fibre of $\pi:V(g,n)\to T(g,n)$ is an immersion?
+By [14, Theorem 1] there exists a holomorphic function $f:V(g,n)\to\mathbb{C}$ without critical points.
+The problem is to find a function $f$ such that $\ker df_{z}$ is transverse to $\ker d\pi_{z}$ at every point $z\in V(g,n)$.
+Note that [16, Corollary 8.3] gives a smooth function $f:V(g,n)\to\mathbb{C}$ whose restriction to any fibre of $\pi:V(g,n)\to T(g,n)$ is a holomorphic immersion.
+Problem 7 is related to the question whether a holomorphic 1-form $\theta$ in Corollary 6 (b) can be made exact on every fibre of $\pi$ by multiplying it with a suitably chosen nowhere vanishing holomorphic function on $V(g,n)$.
+However, this is not the only problem.
+Since the submersion $\pi:V(g,n)\to T(g,n)$ does not admit a holomorphic section when $g\geq 3$, there is no natural way of choosing the initial point for computing the fibrewise integrals of $\theta$, which would give a holomorphic family of immersions on the fibres.
+In [16] the Oka principle was established for families of maps from very general families of open Riemann surfaces $\{(M,J_{t})\}_{t\in T}$ to any Oka manifold $Y$, where $M$ is a smooth open surface and $J_{t}$ are complex structures of some local Hölder class on $M$ depending continuously or smoothly on the parameter $t\in T$ in a suitable topological space.
 The Riemann surfaces in such families need not belong to the same Teichmüller space.
 For example, a punctured Riemann surface can be a member of a family in which the punctures develop into boundary curves and vice versa.
-The $J_{t}$-holomorphic maps $F_{t}:(M,J_{t})\to Y$ $(t\in T)$ furnished by [11, Theorem 1.6] are merely continuous or smooth in the parameter.
+The $J_{t}$-holomorphic maps $F_{t}:(M,J_{t})\to Y$ $(t\in T)$ furnished by [16, Theorem 1.6] are merely continuous or smooth in the parameter.
 In [18] these results were extended to maps from tame families of Stein manifolds of arbitrary dimension to Oka manifolds.
@@
 This is always the case in a family of open Riemann surfaces but it fails general in higher dimensions, even on affine spaces.
-2. Proof of Theorem 1.1
-In view of the description of the Teichmüller submersion $\pi:V(g,n)\to T(g,n)$ in (1.1), Theorem 1.1 is an immediate consequence of the following more general result with arbitrary Stein manifold as the base.
-Theorem 2.1.
-Assume that $X$ is a Stein manifold, $Z$ is a complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion, and $s_{1},\ldots,s_{n}:X\to Z$ are holomorphic sections with pairwise disjoint images for some $n\geq 1$.
+Another interesting question is whether the Riemann surfaces $M_{t}$ in the universal Teichmüller family $V(g,n)$, $n\geq 1$, admit a representation as a family of conformal minimal surfaces in $\mathbb{R}^{k}$, $k\geq 3$, whose $(1,0)$-derivatives depends holomorphically on $t\in T(g,n)$.
+For background, see [29] and [4].
+By [16, Corollary 8.6] the answer is affirmative with continuous or smooth dependence on parameter.
+It remains an open problem whether minimal surfaces in such families can be chosen to be complete and with finite total curvature.
+Each single surface in the family can be made such by [6].
+In view of the description of the Teichmüller submersion $\pi:V(g,n)\to T(g,n)$, Theorem 1 is an immediate consequence of the following result with arbitrary Stein manifold as the base.
+See also Theorem 9 for a more general result.
+Theorem 8.
+Assume that $X$ is a Stein manifold, $Z$ is a connected complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion, and $s_{1},\ldots,s_{n}:X\to Z$ are holomorphic sections with pairwise disjoint images for some $n\geq 1$.
 Then, the domain $\Omega=Z\setminus\bigcup_{i=1}^{n}s_{i}(X)$ is Stein.
@@
 In our case, $Z$ is not compact unless $X$ is a point.
-Proof.
+Proof of Theorem 8.
 Let $\pi:Z\to X$ be as in the theorem.
-We may assume that $X$ is connected.
 Note that every fibre $Z_{x}=\pi^{-1}(x)$, $x\in X$, is a compact Riemann surface, and the fibres are diffeomorphic but not necessarily biholomorphic to each other.
@@
 Each $H_{i}=s_{i}(X)$ is a closed complex hypersurface in $Z$ whose ideal sheaf is a principal, that is, it is locally near each point of $H_{i}$ generated by a single holomorphic function.
-Recall the following result (see Grauert and Remmert [16, Theorem 5, p.
+Recall the following result (see Grauert and Remmert [21, Theorem 5, p.
 129]): If $Z$ is a Stein space and $H$ is a closed complex analytic hypersurface in $Z$ (of pure codimension one) whose ideal sheaf is a principal ideal sheaf, then $Z\setminus H$ is also Stein.
-If $Z$ is nonsingular (a complex manifold) then the ideal sheaf of any closed complex hypersurface in $Z$ is a principal ideal sheaf (see [16, Chap.
+If $Z$ is nonsingular (a complex manifold) then the ideal sheaf of any closed complex hypersurface in $Z$ is a principal ideal sheaf (see [21, Chap.
 A.3.5]).
@@
 By a theorem of Siu [32], the Stein hypersurface $H=s(X)$ has a basis of open Stein neighbourhoods $U$ in $Z$.
-Since $U\setminus H$ is a Stein manifold by the aforementioned theorem [16, Theorem 5, p.
+Since $U\setminus H$ is a Stein manifold by the aforementioned theorem [21, Theorem 5, p.
 129], it admits a strongly plurisubharmonic exhaustion function $\phi:U\setminus H\to\mathbb{R}_{+}$.
@@
 Since $\phi$ tends to $\infty$ along $H$, there are small smoothly bounded open discs $D\Subset D^{\prime}\Subset M$ with $p_{0}\in D$ such that
-(2.1) $\inf_{p\in bD}\phi(x_{0},p)>\max_{p\in bD^{\prime}}\phi(x_{0},p).$
+(2) $\inf_{p\in bD}\phi(x_{0},p)>\max_{p\in bD^{\prime}}\phi(x_{0},p).$
 The set $O=M\setminus D$ is a compact bordered Riemann surface with smooth boundary $bO=bD$, endowed with the complex structure inherited by the identification $M\cong Z_{x_{0}}=\pi^{-1}(x_{0})$.
 Note that $bD^{\prime}$ is contained in the interior of $O$.
-It follows from (2.1) and standard results that there is a smooth strongly subharmonic function $u_{0}:O\to\mathbb{R}_{+}$ such that
+It follows from (2) and standard results that there is a smooth strongly subharmonic function $u_{0}:O\to\mathbb{R}_{+}$ such that
 $u_{0}<\phi(x_{0},\cdotp)$ on $bO=bD$ and $u_{0}>\phi(x_{0},\cdotp)$ on $bD^{\prime}$.
@@
 Note that $\rho_{0}$ is well defined, piecewise smooth, strongly subharmonic on each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X_{0})$, and it agrees with $\phi$ on $(X_{0}\times D)\setminus H$.
-By using the regularised maximum in the definition of $\rho_{0}$ (see [9, Eq.
+By using the regularised maximum in the definition of $\rho_{0}$ (see [15, Eq.
 (3.1), p.
@@
 Hence, if $\tau$ is a strongly plurisubharmonic exhaustion function on $X$ and the function $h:\mathbb{R}\to\mathbb{R}$ is chosen such that $h^{\prime\prime}\geq 0$ and $h^{\prime}$ grows sufficiently fast, then $dd^{c}(h\circ u)$ also grows as fast as desired.
-This completes the proof of Theorem 2.1, and it also proves Theorem 1.1. ∎
-A minor modification of the proof of Theorem 2.1 gives the following more general result.
-Theorem 2.2.
-Assume that $X$ is a connected Stein manifold, $Z$ is a complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion, and $H$ is a closed complex subvariety of $Z$ of pure codimension one which does not contain any fibre of $\pi$.
+This completes the proof of Theorem 8, and it also proves Theorem 1. ∎
+A minor modification of the proof of Theorem 8 gives the following more general result.
+Theorem 9.
+Assume that $X$ is a Stein manifold, $Z$ is a connected complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion, and $H$ is a closed complex subvariety of $Z$ of pure codimension one which does not contain any fibre of $\pi$.
 Then, the domain $\Omega=Z\setminus H$ is Stein.
-Sketch of proof.
-The restricted projection $\pi:H\to X$ is a proper holomorphic map.
+Proof.
+The restricted projection $\pi:H\to X$ is a proper holomorphic map whose fibres are compact Riemann surfaces.
 Since $H$ is of pure codimension one and does not contain any fibre of $\pi$, $\pi:H\to X$ is a finite holomorphic map (a branched holomorphic cover).
-The proper mapping theorem of Remmert [23] (see also Chirka [6, p.
-29]) implies that $\pi(H)$ is a closed complex subvariety of $X$ of pure dimension $n$, hence $\pi(H)=X$ since $X$ is connected, and $H$ is Stein by [16, Theorem 1 (d), p.
+The proper mapping theorem of Remmert [30] (see also Chirka [9, p.
+29]) implies that $\pi(H)$ is a closed complex subvariety of $X$ of pure dimension $n$, hence $\pi(H)=X$ since $X$ is connected, and $H$ is Stein by [21, Theorem 1 (d), p.
 125].
 By Siu’s theorem [32], $H$ admits a basis of open Stein neighbourhoods $U\subset Z$.
-By the argument in the proof of Theorem 2.1, the ideal sheaf of $H$ is a principal ideal sheaf, so $U\setminus H$ is Stein for any Stein neighbourhood $U$ of $H$ by [16, Theorem 5, p.
+By the argument in the proof of Theorem 8, the ideal sheaf of $H$ is a principal ideal sheaf, so $U\setminus H$ is Stein for any Stein neighbourhood $U$ of $H$ by [21, Theorem 5, p.
 129].
-The proof can now be completed by a similar argument as in the proof of Theorem 2.1, and we leave further details to the reader.
+The proof can now be completed by a similar argument as in the proof of Theorem 8, and we leave further details to the reader.
 ∎
@@
 Research was supported by the European Union (ERC Advanced grant HPDR, 101053085) and grants P1-0291 and N1-0237 from ARIS, Republic of Slovenia.
-I wish to thank Finnur Lárusson for posing the question answered by Theorem 1.1 in a private communication.
+I wish to thank Finnur Lárusson for posing the question answered by Theorem 1 in a private communication.

### v4

@@
 Abstract.
-We show that for any pair of integers $g\geq 0$ and $n\geq 1$ the universal family of $n$-punctured compact Riemann surfaces of genus $g$ is a Stein manifold.
+We show that the universal Teichmüller family of $n$-punctured compact Riemann surfaces of genus $g$ is a Stein manifold for any $n>0$.
 We describe its basic function theoretic properties and pose several challenging questions.
+We show in particular that the space of fibrewise algebraic functions on the universal family is dense in the space of holomorphic functions, and there is a fibrewise algebraic map of the universal family in a Euclidean space which is an embedding over any given relatively compact domain in the Teichmüller space.
+We also obtain a relative Oka principle for holomorphic fibrewise algebraic maps of the universal family to any flexible algebraic manifold.
 Key words and phrases:
@@
 Primary 32G15; secondary 32Q28, 32Q56
+1. Introduction
 […]
-(1) $V(g,n)=\widehat{V}(g,n)\setminus\bigcup_{i=1}^{n}s_{i}(T(g,n))$
+(1.1) $V(g,n)=\widehat{V}(g,n)\setminus\bigcup_{i=1}^{n}s_{i}(T(g,n))$
 of $\widehat{V}(g,n)$ is the universal family of $n$-punctured compact Riemann surfaces of genus $g$.
-If $2g+n\geq 3$ then the Teichmüller family $\pi:V(g,n)\to T(g,n)$ is the universal object in the complex analytic category of topologically marked holomorphically varying families of $n$-punctured genus $g$ Riemann surfaces (see [28, Theorem 5.4.3]).
-In this paper we prove the following result.
-Theorem 1.
-If $n\geq 1$ then the universal family $V(g,n)$ is a Stein manifold.
-This is a special case of Theorems 8 and 9 given in the sequel.
-The theorem is trivial if $g=0$ and $n\in\{1,2,3\}$ since $T(0,n)$ is then a singleton and $V(0,n)$ equals $\mathbb{C}$, $\mathbb{C}^{*}=\mathbb{C}\setminus\{0\}$, and $\mathbb{C}\setminus\{0,1\}$, respectively.
-If $2g+n\geq 3$ then $T(g,n)$ is biholomorphic to a bounded topologically contractible Stein domain in $\mathbb{C}^{3g-3+n}$ [28, p.
-161], but it cannot be holomorphically realised as a convex domain in $\mathbb{C}^{3g-3+n}$ if $g\geq 2$ (see Marković [27]).
-The $n$ holomorphic sections $s_{1},\ldots,s_{n}$ of the submersion $\pi:\widehat{V}(g,n)\to T(g,n)$ in (1), corresponding to the punctures of $V(g,n)$, are called canonical sections.
-The problem of describing the global holomorphic sections of this fibration was first considered by Hubbard, who showed in [25] that $V(g,0)\to T(g,0)$ has no holomorphic sections if $g\geq 3$ and six sections if $g=2$.
-See also [26].
-Earle and Kra proved in [10, p.
-50] that for $n\geq 1$, $\pi:\widehat{V}(g,n)\to T(g,n)$ has exactly $n$ sections if $g\geq 3$ and exactly $2n+6$ sections if $g=2$.
-Theorem 1 has interesting consequences for function theory and complex geometry.
-By classical results of Remmert, Bishop, and Narasiman (see [15, Theorem 2.4.1] and the references therein), the Stein manifold $V(g,n)$ for $n\geq 1$ admits a proper holomorphic embedding in $\mathbb{C}^{N}$ with $N=2\dim V(g,n)+1$, which equals 6g6+2n+16g-6+2n+1 if $2g+n\geq 3$.
-When $\dim V(g,n)\geq 2$, it also embeds properly holomorphically in $\mathbb{C}^{N}$ with $N=\big[\frac{3\dim V(g,n)}{2}\big]+1$ (see Eliashberg and Gromov [11] and Schürmann [31], or the exposition in [15, Secs.
+If $2g+n\geq 3$ then the Teichmüller family $\pi:V(g,n)\to T(g,n)$ is the universal object in the complex analytic category of topologically marked holomorphically varying families of $n$-punctured genus $g$ Riemann surfaces (see [39, Theorem 5.4.3]).
+In this paper, we establish several function theoretic properties of the universal family $V(g,n)$. Our first main result is the following.
+Theorem 1.1.
+The Teichmüller family $V(g,n)$ for $n\geq 1$ is a Stein manifold.
+This is a special case of Theorems 2.1 and 2.2 given in the sequel.
+The result is obvious if $g=0$ and $n\in\{1,2,3\}$ since $T(0,n)$ is then a singleton and $V(0,n)$ is biholomorphic to $\mathbb{C}$, $\mathbb{C}^{*}=\mathbb{C}\setminus\{0\}$, and $\mathbb{C}\setminus\{0,1\}$, respectively.
+If $2g+n\geq 3$ then $T(g,n)$ is biholomorphic to a bounded topologically contractible Stein domain in $\mathbb{C}^{3g-3+n}$ [39, p.
+161], but it cannot be holomorphically realised as a convex domain in $\mathbb{C}^{3g-3+n}$ if $g\geq 2$ since the Kobayashi metric and the Carathéodory metric on it differ (see Marković [38]), while they agree on a convex domain (see Lempert [37]).
+Note also that $\widehat{V}(g,n)$ is holomorphically convex and all holomorphic functions on it come from the base $T(g,n)$.
+We provide two proofs of Theorem 1.1.
+The first more elementary one in Sect.
+2 shows that $V(g,n)$ admits a strongly plurisubharmonic exhaustion function, hence is Stein by a theorem of Grauert [26].
+The second one follows from Theorem 3.1 which shows that holomorphic fibrewise algebraic functions on $V(g,n)$ separate points and provide holomorphic convexity.
+The proof uses Grauert’s theorem on coherence of direct images of coherent analytic sheaves and other techniques of complex geometry.
+Theorem 1.1 has several interesting consequences.
+By classical results of Remmert, Bishop, and Narasiman (see [19, Theorem 2.4.1] and the references therein), $V(g,n)$ for $n\geq 1$ admits a proper holomorphic embedding in $\mathbb{C}^{N}$ with $N=2\dim V(g,n)+1$, which equals 6g6+2n+16g-6+2n+1 if $2g+n\geq 3$.
+When $\dim V(g,n)\geq 2$, it also embeds properly holomorphically in $\mathbb{C}^{N}$ with $N=\big[\frac{3\dim V(g,n)}{2}\big]+1$ (see Eliashberg and Gromov [16] and Schürmann [42], or the exposition in [19, Secs.
 9.3–9.4]).
-Since every fibre of $V(g,n)$ is biholomorphic to a closed affine algebraic curve in $\mathbb{C}^{3}$, the following question is natural.
-Problem 2.
-Does $V(g,n)$ for $n\geq 1$ admit a proper holomorphic embedding in some $\mathbb{C}^{N}$ which is algebraic on every fibre?
-Except in the trivial cases, $T(g,n)$ is not affine algebraic, and it is not reasonable to expect that $V(g,n)$ is affine algebraic.
-Since $T(g,n)$ is contractible, the holomorphic submersion $\pi:V(g,n)\to T(g,n)$ is smoothly trivial and the inclusion of any fibre of $\pi$ in $V(g,n)$ is a homotopy equivalence.
-It follows that for any manifold $Y$, the restriction of a continuous map $V(g,n)\to Y$ to any fibre of $V(g,n)$ lies in the same homotopy class of maps $M\to Y$, where $M$ is the underlying $n$-punctured smooth compact surface of genus $g$.
-The following is an immediate corollary to this observation, Theorem 1, and the main result of Oka theory [15, Theorem 5.4.4].
-For Oka manifolds, see [15, Chap.
-5] and the surveys [13, 12, 17].
-Corollary 3.
+Since the Teichmüller space $T(g,n)$ is contractible, the submersion $\pi:V(g,n)\to T(g,n)$ is smoothly trivial and the inclusion of any fibre of $\pi$ in $V(g,n)$ is a homotopy equivalence.
+It follows that for any manifold $Y$, the restriction of a continuous map $V(g,n)\to Y$ to any fibre of $V(g,n)$ lies in the same homotopy class.
+The following is a corollary to this observation, Theorem 1.1, and the main result of Oka theory [19, Theorem 5.4.4].
+(See also the surveys [20, 21, 23].)
+Corollary 1.2.
 Let $\pi:V(g,n)\to T(g,n)$ be as above, $n\geq 1$, and let $Y$ be an Oka manifold.
@@
 The last statement follows by taking the Oka manifold $Y=\mathbb{C}^{*}$.
-Corollary 4.
-If $n\geq 1$, $(g,n)\neq(0,1)$, and $Y$ is an Oka manifold which is not simply connected, there is a holomorphic map $V(g,n)\to Y$ which is nonconstant on every fibre.
+The result obviously holds for g=0,n=1g=0,\ n=1 since $V(0,1)=\mathbb{C}$.
+Corollary 1.3.
+If $n\geq 1$ and $Y$ is an Oka manifold which is not simply connected, there is a holomorphic map $V(g,n)\to Y$ which is nonconstant on every fibre.
 In particular, $V(g,n)$ admits a nowhere vanishing holomorphic function which is nonconstant on every fibre.
-Another way to obtain fibrewise nonconstant holomorphic maps from $V(g,n)$ to an Oka manifold is to inductively use the Oka property with approximation on compact holomorphically convex subsets of the Stein manifold $V(g,n)$; see [15, Theorem 5.4.4].
+Another way to obtain fibrewise nonconstant holomorphic maps from $V(g,n)$ to an Oka manifold is to inductively use the Oka property with approximation on compact holomorphically convex subsets of the Stein manifold $V(g,n)$; see [19, Theorem 5.4.4].
 The Oka principle can be used to obtain many further properties of holomorphic universal families $V(g,n)\to Y$ in any Oka manifold.
+Since every fibre of $V(g,n)$ is an affine algebraic curve, it is natural to expect that $V(g,n)$ admits holomorphic functions which are algebraic on every fibre.
+This is indeed the case.
+We show in Theorem 3.4 that the algebra $\mathscr{A}(V(g,n))$ of fibrewise algebraic functions on $V(g,n)$ is dense in the algebra $\mathscr{O}(V(g,n))$ of holomorphic functions in the compact-open topology.
+Furthermore, functions in $\mathscr{A}(V(g,n))$ provide a proper embedding of $V(g,n)$ in a Euclidean space over any relatively compact domain in $T(g,n)$ (see Theorem 3.1), and fibrewise algebraic maps of $V(g,n)$ to any flexible algebraic manifold satisfy a relative Oka principle (see Theorem 4.1).
+The following problem remains open.
+Problem 1.4.
+Does $V(g,n)$ for $n\geq 1$ admit a proper holomorphic embedding in some $\mathbb{C}^{N}$ which is algebraic on every fibre?
 The fact that $V(g,n)$ for $n\geq 1$ is homotopy equivalent to a bouquet of circles implies that every complex vector bundle on $V(g,n)$ is topologically trivial.
-Since $V(g,n)$ is Stein, the Oka–Grauert principle ([19] or [15, Theorem 3.2.1]) implies the following.
-Proposition 5.
+Since $V(g,n)$ is Stein, the Oka–Grauert principle (see Grauert [25] or [19, Theorem 3.2.1]) implies the following.
+Proposition 1.5.
 Every holomorphic vector bundle on $V(g,n)$ for $n\geq 1$ is holomorphically trivial.
-Corollary 6.
+This fails in the algebraic category, even on a single fibre. In fact, the tangent bundle of a finitely punctured compact Riemann surface is not algebraically trivial in general.
+Corollary 1.6.
 Assume that $g\geq 0$ and $n\geq 1$.
@@
 Proof.
-Part (a) follows by applying Proposition 5 to the holomorphic line bundle $\ker d\pi\to V(g,n)$, the vertical tangent bundle of the holomorphic submersion $\pi:V(g,n)\to T(g,n)$.
+Part (a) follows by applying Proposition 1.5 to the holomorphic line bundle $\ker d\pi\to V(g,n)$, the vertical tangent bundle of the holomorphic submersion $\pi:V(g,n)\to T(g,n)$.
 To see (b), consider the following short exact sequence of vector bundles over $V(g,n)$:
@@
 ∎
-By the Gunning–Narasimhan theorem [23], every open Riemann surface $M$ admits a holomorphic immersion $f:M\to\mathbb{C}$. In view of Corollary 6 (b), the following is a natural question.
-Problem 7.
-Let $n\geq 1$. Does there exist a holomorphic function $f:V(g,n)\to\mathbb{C}$ whose restriction to any fibre of $\pi:V(g,n)\to T(g,n)$ is an immersion?
-By [14, Theorem 1] there exists a holomorphic function $f:V(g,n)\to\mathbb{C}$ without critical points.
-The problem is to find a function $f$ such that $\ker df_{z}$ is transverse to $\ker d\pi_{z}$ at every point $z\in V(g,n)$.
-Note that [16, Corollary 8.3] gives a smooth function $f:V(g,n)\to\mathbb{C}$ whose restriction to any fibre of $\pi:V(g,n)\to T(g,n)$ is a holomorphic immersion.
-Problem 7 is related to the question whether a holomorphic 1-form $\theta$ in Corollary 6 (b) can be made exact on every fibre of $\pi$ by multiplying it with a suitably chosen nowhere vanishing holomorphic function on $V(g,n)$.
+By the Gunning–Narasimhan theorem [31], every open Riemann surface $M$ admits a holomorphic immersion $f:M\to\mathbb{C}$. In view of Corollary 1.6 (b), the following is a natural question.
+Problem 1.7.
+Let $n\geq 1$. Is there a holomorphic function $f:V(g,n)\to\mathbb{C}$ whose restriction to every fibre of $V(g,n)$ is an immersion?
+By [17, Theorem 1] there exists a holomorphic function $f:V(g,n)\to\mathbb{C}$ without critical points.
+The problem is to find $f$ such that $\ker df_{z}$ is transverse to $\ker d\pi_{z}$ at every point $z\in V(g,n)$.
+Note that [22, Corollary 8.3] gives a smooth function $f:V(g,n)\to\mathbb{C}$ whose restriction to any fibre is a holomorphic immersion.
+Problem 1.7 is related to the question whether a holomorphic 1-form $\theta$ in Corollary 1.6 (b) can be made exact on every fibre of $\pi$ by multiplying it with a suitably chosen nowhere vanishing holomorphic function on $V(g,n)$.
 However, this is not the only problem.
-Since the submersion $\pi:V(g,n)\to T(g,n)$ does not admit a holomorphic section when $g\geq 3$, there is no natural way of choosing the initial point for computing the fibrewise integrals of $\theta$, which would give a holomorphic family of immersions on the fibres.
-In [16] the Oka principle was established for families of maps from very general families of open Riemann surfaces $\{(M,J_{t})\}_{t\in T}$ to any Oka manifold $Y$, where $M$ is a smooth open surface and $J_{t}$ are complex structures of some local Hölder class on $M$ depending continuously or smoothly on the parameter $t\in T$ in a suitable topological space.
+Since the Teichmüller submersion $\pi:V(g,n)\to T(g,n)$ does not admit a holomorphic section when $g\geq 3$ (see Hubbard [34, 35] and Earle and Kra [15, p.
+50] for a precise description of sections of $\widehat{V}(g,n)$ and $V(g,n)$), there is no natural way of choosing the initial point for computing the fibrewise integrals of $\theta$, which would give a holomorphic family of immersions on the fibres.
+Another interesting question is whether the Riemann surfaces $M_{t}$ in the Teichmüller family $V(g,n)$, $n\geq 1$, admit a representation as a family of conformal minimal surfaces in $\mathbb{R}^{k}$, $k\geq 3$, whose $(1,0)$-derivatives depends holomorphically on $t\in T(g,n)$.
+For background, see [40] and [4].
+By [22, Corollary 8.6] the answer is affirmative with continuous or smooth dependence on parameter.
+It remains an open problem whether minimal surfaces in such families can be chosen to be complete and with finite total curvature.
+Each single surface in the family can be made such by [7, 6].
+It is natural to wonder whether the results of this paper extend to infinite dimensional Teichmüller families.
+The analogous problem with continuous or smooth dependence of the complex structures and the holomorphic maps on the parameter was studied in [22] for very general families of open Riemann surfaces $\{(M,J_{t})\}_{t\in T}$, where $M$ is a smooth open surface and $J_{t}$ are complex structures of some local Hölder class on $M$ depending continuously or smoothly on the parameter $t$ in a topological space $T$.
 The Riemann surfaces in such families need not belong to the same Teichmüller space.
 For example, a punctured Riemann surface can be a member of a family in which the punctures develop into boundary curves and vice versa.
-The $J_{t}$-holomorphic maps $F_{t}:(M,J_{t})\to Y$ $(t\in T)$ furnished by [16, Theorem 1.6] are merely continuous or smooth in the parameter.
-In [18] these results were extended to maps from tame families of Stein manifolds of arbitrary dimension to Oka manifolds.
-Here, a continuous family $\{J_{t}\}_{t\in T}$ of Stein structures on a smooth manifold $X$ said to be tame if the holomorphically convex hulls of any compact subset $K\subset X$ are upper semicontinuous with respect to the parameter $t\in T$.
-This is always the case in a family of open Riemann surfaces but it fails general in higher dimensions, even on affine spaces.
-Another interesting question is whether the Riemann surfaces $M_{t}$ in the universal Teichmüller family $V(g,n)$, $n\geq 1$, admit a representation as a family of conformal minimal surfaces in $\mathbb{R}^{k}$, $k\geq 3$, whose $(1,0)$-derivatives depends holomorphically on $t\in T(g,n)$.
-For background, see [29] and [4].
-By [16, Corollary 8.6] the answer is affirmative with continuous or smooth dependence on parameter.
-It remains an open problem whether minimal surfaces in such families can be chosen to be complete and with finite total curvature.
-Each single surface in the family can be made such by [6].
-In view of the description of the Teichmüller submersion $\pi:V(g,n)\to T(g,n)$, Theorem 1 is an immediate consequence of the following result with arbitrary Stein manifold as the base.
-See also Theorem 9 for a more general result.
-Theorem 8.
-Assume that $X$ is a Stein manifold, $Z$ is a connected complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion, and $s_{1},\ldots,s_{n}:X\to Z$ are holomorphic sections with pairwise disjoint images for some $n\geq 1$.
+Under mild assumptions on the family $\{J_{t}\}_{t\in T}$, $J_{t}$-holomorphic functions on $(M,J_{t})$ satisfy the Runge approximation theorem with continuous or smooth dependence on $t\in T$ [22, Theorem 1.1].
+For a class of parameter spaces including finite CW complexes we also have the Oka principle for continuous or smooth families of holomorphic maps from $(M,J_{t})$ to any Oka manifold [22, Theorem 1.6].
+These results were extended in [24] to maps from tame families of Stein manifolds of arbitrary dimension to Oka manifolds.
+2. Proof of Theorem 1.1
+In view of the description of the Teichmüller submersion $\pi:V(g,n)\to T(g,n)$ (1.1), Theorem 1.1 is an immediate consequence of the following result with arbitrary Stein manifold as the base.
+Theorem 2.1.
+Assume that $X$ is a Stein manifold, $Z$ is a complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion with connected fibres, and $s_{1},\ldots,s_{n}:X\to Z$ $(n\geq 1)$ are holomorphic sections with pairwise disjoint images.
 Then, the domain $\Omega=Z\setminus\bigcup_{i=1}^{n}s_{i}(X)$ is Stein.
-The conclusion fails if the fibres of $\pi$ have complex dimension >1>1 or the sections $s_{i}$ are not holomorphic.
-In such case, the domain $\Omega$ in the theorem fails to be locally pseudoconvex at some boundary point $s_{i}(x)$, $x\in X$.
+The assumption that the sections $s_{1},\ldots,s_{n}$ have pairwise disjoint images is inessential and is imposed only for convenience of the proof; see Theorem 2.2 for a more general result.
+The conclusion fails if the fibres of $\pi$ have complex dimension >1>1, or if the sections $s_{i}$ are not holomorphic.
+In such a case, the domain $\Omega$ in the theorem fails to be locally pseudoconvex at some boundary point $s_{i}(x)$, $x\in X$.
 Note that Stein complements of smooth complex hypersurfaces in compact Kähler manifolds have recently been studied by Höring and Peternell [32] where the reader can find references to earlier works.
 In our case, $Z$ is not compact unless $X$ is a point.
-Proof of Theorem 8.
+Proof of Theorem 2.1.
 Let $\pi:Z\to X$ be as in the theorem.
-Note that every fibre $Z_{x}=\pi^{-1}(x)$, $x\in X$, is a compact Riemann surface, and the fibres are diffeomorphic but not necessarily biholomorphic to each other.
+Note that every fibre $Z_{x}=\pi^{-1}(x)$, $x\in X$, is a compact connected Riemann surface, and the fibres are diffeomorphic but not necessarily biholomorphic to each other.
 Hence, $\{Z_{x}\}_{x\in X}$ is a holomorphic family of compact Riemann surfaces and $\Omega_{x}=Z_{x}\setminus\bigcup_{i=1}^{n}s_{i}(x)$ $(x\in X)$ is a holomorphic family of $n$-punctured Riemann surfaces.
-Each $H_{i}=s_{i}(X)$ is a closed complex hypersurface in $Z$ whose ideal sheaf is a principal, that is, it is locally near each point of $H_{i}$ generated by a single holomorphic function.
-Recall the following result (see Grauert and Remmert [21, Theorem 5, p.
+Each $H_{i}=s_{i}(X)$ is a closed complex hypersurface in $Z$ whose ideal sheaf is principal, that is, it is locally near each point of $H_{i}$ generated by a single holomorphic function.
+Recall the following result (see Grauert and Remmert [27, Theorem 5, p.
 129]): If $Z$ is a Stein space and $H$ is a closed complex analytic hypersurface in $Z$ (of pure codimension one) whose ideal sheaf is a principal ideal sheaf, then $Z\setminus H$ is also Stein.
-If $Z$ is nonsingular (a complex manifold) then the ideal sheaf of any closed complex hypersurface in $Z$ is a principal ideal sheaf (see [21, Chap.
+If $Z$ is nonsingular then the ideal sheaf of any closed complex hypersurface in $Z$ is a principal ideal sheaf (see [27, Chap.
 A.3.5]).
-Hence, is suffices to prove the theorem in the case $n=1$, that is, to show that the complement $Z\setminus s(X)$ of a single holomorphic section $s:X\to Z$ is a Stein manifold.
+Hence, it suffices to prove the theorem in the case $n=1$, that is, to show that the complement $Z\setminus s(X)$ of a holomorphic section $s:X\to Z$ is a Stein manifold.
 By a theorem of Siu [44], the Stein hypersurface $H=s(X)$ has a basis of open Stein neighbourhoods $U$ in $Z$.
-Since $U\setminus H$ is a Stein manifold by the aforementioned theorem [21, Theorem 5, p.
+Since $U\setminus H$ is a Stein manifold [27, Theorem 5, p.
 129], it admits a strongly plurisubharmonic exhaustion function $\phi:U\setminus H\to\mathbb{R}_{+}$.
@@
 Since $\phi$ tends to $\infty$ along $H$, there are small smoothly bounded open discs $D\Subset D^{\prime}\Subset M$ with $p_{0}\in D$ such that
-(2) $\inf_{p\in bD}\phi(x_{0},p)>\max_{p\in bD^{\prime}}\phi(x_{0},p).$
+(2.1) $\inf_{p\in bD}\phi(x_{0},p)>\max_{p\in bD^{\prime}}\phi(x_{0},p).$
 The set $O=M\setminus D$ is a compact bordered Riemann surface with smooth boundary $bO=bD$, endowed with the complex structure inherited by the identification $M\cong Z_{x_{0}}=\pi^{-1}(x_{0})$.
 Note that $bD^{\prime}$ is contained in the interior of $O$.
-It follows from (2) and standard results that there is a smooth strongly subharmonic function $u_{0}:O\to\mathbb{R}_{+}$ such that
+It follows from (2.1) and standard results that there is a smooth strongly subharmonic function $u_{0}:O\to\mathbb{R}_{+}$ such that
 $u_{0}<\phi(x_{0},\cdotp)$ on $bO=bD$ and $u_{0}>\phi(x_{0},\cdotp)$ on $bD^{\prime}$.
@@
 Note that $\rho_{0}$ is well defined, piecewise smooth, strongly subharmonic on each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X_{0})$, and it agrees with $\phi$ on $(X_{0}\times D)\setminus H$.
-By using the regularised maximum in the definition of $\rho_{0}$ (see [15, Eq.
+By using the regularised maximum in the definition of $\rho_{0}$ (see [19, Eq.
 (3.1), p.
 69]), we may assume that $\rho_{0}$ is smooth and enjoys the other stated properties.
-This construction gives an open locally finite cover $\{X_{j}\}_{j=1}^{\infty}$ of $X$ with smooth fibre bundle trivialisations $Z|X_{j}\cong X_{j}\times M$, discs $D_{j}\subset M$ such that $s(x)\in D_{j}$ for all $x\in X_{j}$, and smooth functions $\rho_{j}:(Z|X_{j})\setminus H=\pi^{-1}(X_{j})\setminus H\to\mathbb{R}_{+}$ such that $\rho_{j}$ is strongly subharmonic on each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X_{j})$ and it agrees with $\phi$ on $(X_{j}\times D_{j})\setminus H$.
+This construction gives an open locally finite cover $\{X_{j}\}_{j=1}^{\infty}$ of $X$ with smooth fibre bundle trivialisations $Z|X_{j}\cong X_{j}\times M$, discs $D_{j}\subset M$ such that $s(x)\in D_{j}$ for all $x\in X_{j}$, and smooth functions $\rho_{j}:(Z|X_{j})\setminus H\to\mathbb{R}_{+}$ such that $\rho_{j}$ is strongly subharmonic on each fibre $Z_{x}\setminus\{s(x)\}$ $(x\in X_{j})$ and it agrees with $\phi$ on $(X_{j}\times D_{j})\setminus H$.
 Let $\{\chi_{j}\}_{j}$ be a smooth partition of unity on $X$ with compact supports $\mathrm{supp}(\chi_{j})\subset X_{j}$ for each $j$.
@@
 Note that for every compact set $K\subset X$, the set $\pi^{-1}(K)\setminus U_{0}\subset Z\setminus H$ is compact.
-Hence, by choosing a strongly plurisubharmonic exhaustion function $\tau:X\to\mathbb{R}_{+}$ whose Levi form $dd^{c}\tau$ grows fast enough, we can ensure that $\rho+\tau\circ\pi:Z\setminus H\to\mathbb{R}_{+}$ is a strongly plurisubharmonic exhaustion function.
-Indeed, denoting by $J$ the almost complex structure operator on $Z$ and by $d^{c}$ the conjugate differential defined by $(d^{c}\rho)(z,\xi)=-d\rho(z,J\xi)$ for $z\in Z$ and $\xi\in T_{z}Z$, $\rho$ is strongly plurisubharmonic at $z\in Z$ if and only if $(dd^{c}\rho)(z,\xi\wedge J\xi)>0$ for every tangent vector $0\neq\xi\in T_{z}Z$.
-(Up to a positive factor this equals the Laplace of $\rho$ on the 2-plane $\mathrm{span}(\xi,J\xi)\subset T_{z}Z$.) Since the function $\rho$ constructed above is strongly subharmonic on every fibre $Z_{x}\setminus\{s(x)\}$, $x\in X$, we have that
+Hence, by choosing a strongly plurisubharmonic exhaustion function $\tau:X\to\mathbb{R}_{+}$ whose Levi form $dd^{c}\tau$ grows fast enough, we can ensure that
+$\rho+\tau\circ\pi:Z\setminus H\to\mathbb{R}_{+}$
+is a strongly plurisubharmonic exhaustion function, thus proving the theorem. Indeed, denoting by $J$ the almost complex structure operator on $Z$ and by $d^{c}$ the conjugate differential defined by
+$(d^{c}\rho)(z,\xi)=-d\rho(z,J\xi)\ \ \text{for $z\in Z$ and $\xi\in T_{z}Z$},$
+the function $\rho$ is strongly plurisubharmonic at $z\in Z$ if and only if
+$(dd^{c}\rho)(z,\xi\wedge J\xi)>0\ \ \text{for every tangent vector $0\neq\xi\in T_{z}Z$}.$
+(Up to a positive factor this equals the Laplace of $\rho$ on the 2-plane $\mathrm{span}(\xi,J\xi)\subset T_{z}Z$.) Since $\rho$ is strongly subharmonic on every fibre $Z_{x}\setminus\{s(x)\}$, $x\in X$, we have that
 $(dd^{c}\rho)(z,\xi\wedge J\xi)>0\ \ \text{if $z\in Z\setminus H$ and $0\neq\xi\in\ker d\pi_{z}$.}$
-Hence, the eigenvectors of the Levi form $(dd^{c}\rho)(z,\cdotp)$ associated to non-positive eigenvalues lie in a closed cone $C_{z}\subset T_{z}Z$ which intersects $\ker d\pi_{z}$ only in the origin.
-It follows that if $\tau:X\to\mathbb{R}$ is such that $dd^{c}\tau>0$ is sufficiently big on $T_{x}X$, $x=\pi(z)$, then $dd^{c}\rho+dd^{c}(\tau\circ\pi)>0$ on $T_{z}Z$.
-Furthermore, the estimates are uniform on the compact set $\pi^{-1}(K)\setminus U_{0}$, where $U_{0}\subset Z$ is a neighbourhood of $H$ as above such that $dd^{c}\rho>0$ on $U_{0}\setminus H$.
-To see that $\tau$ can be chosen such that $dd^{c}\tau$ grows as fast as desired, note that if $h:\mathbb{R}\to\mathbb{R}$ is a $\mathscr{C}^{2}$ function then for each point $x\in X$ and tangent vector $\xi\in T_{x}X$ we have that
-$dd^{c}(h\circ\tau)(x,\xi\wedge J\xi)=h^{\prime}(\tau(x))\,(dd^{c}\tau)(x,\xi\wedge J\xi)+h^{\prime\prime}(\tau(x))\,\big(|d\tau(x,\xi)|^{2}+|d\tau(x,J\xi|^{2}\big).$
-Hence, if $\tau$ is a strongly plurisubharmonic exhaustion function on $X$ and the function $h:\mathbb{R}\to\mathbb{R}$ is chosen such that $h^{\prime\prime}\geq 0$ and $h^{\prime}$ grows sufficiently fast, then $dd^{c}(h\circ u)$ also grows as fast as desired.
-This completes the proof of Theorem 8, and it also proves Theorem 1. ∎
-A minor modification of the proof of Theorem 8 gives the following more general result.
-Theorem 9.
-Assume that $X$ is a Stein manifold, $Z$ is a connected complex manifold with $\dim Z=\dim X+1$, $\pi:Z\to X$ is a surjective proper holomorphic submersion, and $H$ is a closed complex subvariety of $Z$ of pure codimension one which does not contain any fibre of $\pi$.
+Hence, the eigenvectors of $(dd^{c}\rho)(z,\cdotp)$ associated to non-positive eigenvalues lie in a closed cone $C_{z}\subset T_{z}Z$ which intersects $\ker d\pi_{z}$ only in the origin.
+It follows that if $\tau:X\to\mathbb{R}$ is such that $dd^{c}\tau>0$ is sufficiently big on $T_{x}X$ where $x=\pi(z)$, then $dd^{c}\rho+dd^{c}(\tau\circ\pi)>0$ on $T_{z}Z$.
+Furthermore, the estimates are uniform on the compact set $\pi^{-1}(K)\setminus U_{0}$, where $U_{0}\subset Z$ is a neighbourhood of $H$ such that $dd^{c}\rho>0$ on $U_{0}\setminus H$.
+To see that $\tau$ can be chosen such that $dd^{c}\tau$ grows as fast as desired, note that if $h:\mathbb{R}\to\mathbb{R}$ is a $\mathscr{C}^{2}$ function then for each point $x\in X$ and vector $\xi\in T_{x}X$ we have that
+$dd^{c}(h\circ\tau)(x,\xi\wedge J\xi)=h^{\prime}(\tau(x))\,(dd^{c}\tau)(x,\xi\wedge J\xi)+h^{\prime\prime}(\tau(x))\,\big(|d\tau(x,\xi)|^{2}+|d\tau(x,J\xi)|^{2}\big).$
+Hence, if $\tau$ is a strongly plurisubharmonic exhaustion function on $X$ and the function $h:\mathbb{R}\to\mathbb{R}$ is chosen such that $h^{\prime\prime}\geq 0$ and $h^{\prime}$ grows sufficiently fast, then $dd^{c}(h\circ\tau)$ also grows as fast as desired.
+∎
+A minor modification of the proof of Theorem 2.1 gives the following more general result.
+Theorem 2.2.
+Assume that $X$ is a connected Stein manifold, $Z$ is a complex manifold, $\pi:Z\to X$ is a surjective proper holomorphic submersion with connected one-dimensional fibres, and $H$ is a closed complex subvariety of $Z$ of pure codimension one which does not contain any fibre of $\pi$.
 Then, the domain $\Omega=Z\setminus H$ is Stein.
 Proof.
-The restricted projection $\pi:H\to X$ is a proper holomorphic map whose fibres are compact Riemann surfaces.
-Since $H$ is of pure codimension one and does not contain any fibre of $\pi$, $\pi:H\to X$ is a finite holomorphic map (a branched holomorphic cover).
-The proper mapping theorem of Remmert [30] (see also Chirka [9, p.
-29]) implies that $\pi(H)$ is a closed complex subvariety of $X$ of pure dimension $n$, hence $\pi(H)=X$ since $X$ is connected, and $H$ is Stein by [21, Theorem 1 (d), p.
-125].
-By Siu’s theorem [32], $H$ admits a basis of open Stein neighbourhoods $U\subset Z$.
-By the argument in the proof of Theorem 8, the ideal sheaf of $H$ is a principal ideal sheaf, so $U\setminus H$ is Stein for any Stein neighbourhood $U$ of $H$ by [21, Theorem 5, p.
+The hypotheses imply that $\pi|_{H}:H\to X$ is a finite holomorphic map (a branched holomorphic cover).
+The proper mapping theorem of Remmert [41] (see also [12, p.
+29]) implies that $\pi(H)$ is a closed complex subvariety of $X$ of pure dimension $\dim X$, hence $\pi(H)=X$ since $X$ is connected.
+By [27, Theorem 1 (d), p.
+125], the subvariety $H$ is Stein.
+A theorem of Siu [44] implies that $H$ has a basis of open Stein neighbourhoods $U\subset Z$.
+By the argument in the proof of Theorem 2.1, the ideal sheaf of $H$ is a principal ideal sheaf, so the manifold $U\setminus H$ is Stein for any Stein neighbourhood $U\subset Z$ of $H$ [27, Theorem 5, p.
 129].
-The proof can now be completed by a similar argument as in the proof of Theorem 8, and we leave further details to the reader.
+The proof can now be completed by a similar argument as in the proof of Theorem 2.1, and we leave the details to the reader.
 ∎
+Remark 2.3.
+Theorem 2.2 still holds if the fibres of the projection $\pi:Z\to X$ are not necessarily connected provided that the subvariety $H$ intersects every connected component of each fibre $Z_{x}=\pi^{-1}(x),\ x\in X$, but it does not contain any such component.
+The result also holds if $X$ is a Stein space, $Z$ is a complex space, $H\subset Z$ is a complex hypersurface whose ideal sheaf is principal, and the other conditions on the submersion $\pi:Z\to X$ and $H$ are satisfied.
+3. Fibrewise algebraic functions on $V(g,n)$
+In this section, we show that the universal family $V(g,n)$ of $n$-punctured compact Riemann surfaces of genus $g$ for $n\geq 1$ has a large algebra of holomorphic functions which are algebraic on every fibre of the Teichmüller projection $\pi:V(g,n)\to T(g,n)$.
+Indeed, such functions are dense in the algebra of holomorphic functions in the compact-open topology; see Theorem 3.4.
+We shall consider the more general situation given in Theorem 2.2.
+Thus, we assume that $X$ is a connected Stein manifold, $Z$ is a complex manifold, $\pi:Z\to X$ is a surjective proper holomorphic submersion with connected one-dimensional fibres, and $H$ is a closed complex subvariety of $Z$ of pure codimension one which does not contain any fibre of $\pi$.
+The domain $\Omega=Z\setminus H$ is Stein by Theorem 2.2, and every fibre $\Omega_{x}$ $(x\in X)$ is a finitely punctured compact Riemann surface, hence affine.
+Let $\mathscr{A}(\Omega)$ denote the $\mathbb{C}$-algebra of holomorphic functions on $\Omega$ which are algebraic on every fibre $\Omega_{x}$, $x\in X$.
+Clearly, $\mathscr{A}(\Omega)$ is a subalgebra of the algebra $\mathscr{O}(\Omega)$ of holomorphic functions.
+Theorem 3.1.
+(Assumptions as above.) Given a relatively compact domain $U\Subset X$, there exist finitely many functions $f_{1},\ldots,f_{n}\in\mathscr{A}(\Omega)$ such that the map $F:\Omega\to X\times\mathbb{C}^{n}$ given by
+(3.1) $F(z)=(\pi(z),f_{1}(z),\ldots,f_{n}(z)),\quad z\in\Omega$
+induces a proper embedding $\Omega_{U}:=\Omega\cap\pi^{-1}(U)\to U\times\mathbb{C}^{n}$.
+A map $F$ satisfying the conclusion of theorem is called a proper embedding over $U$.
+The same result holds if $X$ is a Stein space, $Z$ is a complex space, and the other conditions on the submersion $\pi:Z\to X$ and the hypersurface $H\subset Z$ hold; in particular, the ideal sheaf of $H$ is principal.
+Theorem 3.1 clearly implies that holomorphic fibrewise algebraic functions on $\Omega$ separate points and provide holomorphic convexity, thus providing another proof that $\Omega$ (and in particular the universal family $V(g,n)$) is a Stein manifold.
+Proof.
+Let $L\to Z$ denote the holomorphic line bundle determined by the divisor $H\subset X$, and let $\sigma_{0}:Z\to L$ be a holomorphic section whose zero divisor equals $H$.
+(The existence of such a section is tautological from the construction of $L$.) Since the fibre $Z_{x}=\pi^{-1}(x)$ is a projective curve for every $x\in X$, Serre’s GAGA principle [43] implies that the restricted line bundle $L_{x}:=L|Z_{x}\to Z_{x}$ is algebraic, and every holomorphic section of $L_{x}$ over $Z_{x}$ is algebraic.
+It follows that for any holomorphic section $\sigma:Z\to L$, the quotient $f=\sigma/\sigma_{0}$ is a well defined function in $\mathscr{A}(\Omega)$.
+The restriction of $f$ to $\Omega_{x}$ is algebraic and has an effective pole at every end of $\Omega_{x}$ (that is, a point of $H_{x}=H\cap Z_{x}$) at which $\sigma$ does not vanish.
+Note that the line bundle $L_{x}\to Z_{x}$ is associated to the effective divisor supported on $H_{x}$ in which the multiplicity of a point $z\in H_{x}$ is the intersection number of $H$ with $Z_{x}$ at $z$.
+In particular, $L_{x}$ is an ample line bundle on $Z_{x}$ for every $x\in X$.
+Hence, some tensor power $L_{x}^{\otimes d}\to Z_{x}$ for $d=d(x)>0$ is very ample, so it admits finitely many holomorphic sections $\sigma_{1,x},\cdots,\sigma_{n,x}:Z_{x}\to L_{x}$ such that the map
+(3.2) $Z_{x}\ni z\longmapsto\big[\sigma_{0,x}^{d}(z):\sigma_{1,x}(z):\cdots:\sigma_{n,x}(z)\big]\in\mathbb{CP}^{n}$
+is a holomorphic embedding.
+Here, $\sigma_{0}^{d}$ denotes the $d$-th power of $\sigma_{0}$, a section of the line bundle $L^{\otimes d}\to Z$.
+Note that $\sigma_{i,x}/\sigma_{0,x}^{d}$ is a regular algebraic function on $\Omega_{x}$ for every $i=1,\ldots,n$.
+To prove the theorem, we will show that for $d=d(x)>0$ big enough, every holomorphic section $\sigma_{x}:Z_{x}\to L_{x}^{\otimes d}$ extends to a holomorphic section $\sigma:Z\to L^{\otimes d}$.
+Assume for a moment that this holds true.
+Extending all section $\sigma_{i,x}$ in (3.2) for $i=1,\ldots,n$ to $Z$, the map in (3.2) is an embedding for any base point in an open neighbourhood $U_{1}\subset X$ of $x$.
+Hence, the map $F$ of the form (3.1) with the components $f_{i}=\sigma_{i}/\sigma_{0}^{d}\in\mathscr{A}(\Omega)$ is a proper embedding over $U_{1}$.
+Indeed, for any pair of points $x\in U_{1}$ and $z\in H_{x}$, at least one of the sections $\sigma_{i,x}$ $(i=1,\ldots,n)$ does not vanish at $z$.
+Since $\sigma_{0}(z)=0$, it follows that $f_{i}=\sigma_{i}/\sigma_{0}^{d}$ has an effective pole at $z$, which implies properness.
+Repeating the same construction at other points $x\in\overline{U}$ (for different values of $d=d(x)\in\mathbb{N}$) and assembling the resulting functions as components of a map (3.1) yields a proper embedding over $U$ as in the theorem.
+It remains to explain why every holomorphic section $\sigma_{x}:Z_{x}\to L_{x}^{\otimes d}$ for $d>0$ big enough extends to a holomorphic section $\sigma:Z\to L^{\otimes d}$.
+Let $\mathscr{L}\to Z$ denote the sheaf of holomorphic sections of the line bundle $L\to Z$; note that $\mathscr{L}$ is a coherent $\mathscr{O}_{Z}$-analytic sheaf.
+Since the holomorphic projection $\pi:Z\to X$ is surjective and proper, the direct image sheaf $\pi_{*}\mathscr{L}$ is a coherent analytic sheaf on $X$ by Grauert’s direct image theorem [28, p.
+207].
+A section of $\pi_{*}\mathscr{L}$ over an open subset $V\subset X$ is the same thing as a section of $\mathscr{L}$ over $Z_{V}:=\pi^{-1}(V)$.
+The argument given (in the algebraic case) in [45, Tag 0D2M] shows that for $d_{0}=d_{0}(x)>0$ big enough, the evaluation map
+(3.3) $E_{x}:\pi_{*}(\mathscr{L}^{\otimes d})_{x}\to H^{0}(Z_{x},\mathscr{L}_{x}^{\otimes d})$
+is surjective for $d\geq d_{0}$.
+This means that every holomorphic section $\sigma_{x}:Z_{x}\to L_{x}^{\otimes d}$ of the line bundle $L_{x}^{\otimes d}\to Z_{x}$ extends to a holomorphic section $\sigma$ of $L^{\otimes d}|{Z_{V}}\to Z_{V}$ for an open neighbourhood $V$ of $x$.
+Denote by $\sigma]_{x}\in\pi_{*}(\mathscr{L}^{\otimes d})_{x}$ the associated element of the stalk $\pi_{*}(\mathscr{L}^{\otimes d})_{x}$, so $E_{x}([\sigma]_{x})=\sigma_{x}$.
+Since the sheaf $\pi_{*}\mathscr{L}^{\otimes d}$ is $\mathscr{O}_{X}$-coherent, Cartan’s Theorem A [27, p.
+124] shows that $\pi_{*}(\mathscr{L}^{\otimes d})_{x}$ is generated as an $\mathscr{O}_{X,x}$-module by global sections of $\pi_{*}(\mathscr{L}^{\otimes d})$.
+Hence, there are sections $\xi_{1},\ldots,\xi_{m}\in\Gamma(X,\pi_{*}(\mathscr{L}^{\otimes d}))$ and germs of holomorphic functions $g_{1},\ldots,g_{m}\in\mathscr{O}_{X,x}$ such that
+(3.4) $\sigma]_{x}=\sum_{j=1}^{m}g_{j}[\xi_{j}]_{x}\in\pi_{*}(\mathscr{L}^{\otimes d})_{x}.$
+Since $X$ is Stein, there are functions $\tilde{g}_{j}\in\mathscr{O}(X)$ such that $\tilde{g}_{j}(x)=g_{j}(x)$ for $j=1,\ldots,m$.
+Then, $\tilde{\sigma}:=\sum_{j=1}^{m}\tilde{g}_{j}\xi_{j}\in\Gamma(X,\pi_{*}(\mathscr{L}^{\otimes d}))$ is a section whose value at $x$ equals
+$E_{x}(\tilde{\sigma})=\sum_{j=1}^{m}\tilde{g}_{j}(x)E_{x}(\xi_{j})=\sum_{j=1}^{m}g_{j}(x)E_{x}(\xi_{j})=E_{x}([\sigma]_{x})=\sigma_{x}$
+(see (3.4)). This completes the proof. ∎
+Remark 3.2.
+The integer $d_{0}=d_{0}(x)>0$ in the above argument may depend on $x$.
+For this reason, we do not know whether there is a proper embedding $F:\Omega\to X\times\mathbb{C}^{n}$ of class $\mathscr{A}(\Omega)$ and of the form (3.1) for some $n\in\mathbb{N}$.
+However, such an embedding exists and can be chosen fully algebraic if the base $X$ is affine algebraic due to finiteness of the above procedure, using the Zariski topology as opposed to the Hausdorff topology.
+The following is a special case of a lemma due to R.
+van Dobben de Bruyn [13].
+This result will not be used in the paper and is stated only for the sake of comparison.
+Lemma 3.3.
+Assume that $X$ is an affine algebraic variety, $Z$ is an algebraic variety, $\pi:Z\to X$ is a proper algebraic submersion with one dimensional fibres, and $s_{1},\cdots,s_{m}$ for $m\geq 1$ are algebraic sections.
+If all fibres $Z_{x}\setminus\bigcup_{i=1}^{m}s_{i}(x)$ $(x\in X)$ are affine then there exists a proper algebraic morphism $F:\Omega=Z\setminus\bigcup_{i=1}^{m}s_{i}(X)\to X\times\mathbb{C}^{n}$ of the form (3.1) for some $n\in\mathbb{N}$.
+Recall (see e.g. [33]) that a compact set $K$ in a complex manifold $\Omega$ is said to be $\mathscr{O}(\Omega)$-convex if it equals its $\mathscr{O}(\Omega)$-convex hull
+$\widehat{K}_{\mathscr{O}(\Omega)}=\{z\in\Omega:|f(z)|\leq\sup_{K}|f|\ \text{for all}\ f\in\mathscr{O}(\Omega)\}.$
+The following Oka–Weil type approximation theorem for the Stein domain $\Omega=Z\setminus H$ in Theorem 3.1 shows that the algebra $\mathscr{A}(\Omega)$ is dense in $\mathscr{O}(\Omega)$ in the compact-open topology.
+Theorem 3.4.
+Let $\pi:Z\to X$, $H\subset Z$, and $\Omega=Z\setminus H$ be as in Theorem 3.1.
+Given a compact $\mathscr{O}(X)$-convex set $K\subset\Omega$, every holomorphic function on a neighbourhood of $K$ is a uniform limit on $K$ of functions in $\mathscr{A}(\Omega)$.
+This holds in particular for the Teichmüller family $\Omega=V(g,n)$ (1.1) for any $g\geq 0$ and $n\geq 1$.
+Proof.
+Choose a relatively compact Runge domain $U\Subset X$ such that $\pi(K)\subset U$.
+Let $F:\Omega\to X\times\mathbb{C}^{n}$ be a map of class $\mathscr{A}(\Omega)$, furnished by Theorem 3.1 (see (3.1)), such that $F:\Omega_{U}\to U\times\mathbb{C}^{n}$ is a proper embedding.
+Its image $F(\Omega_{U})$ is then a closed complex submanifold of the Stein manifold $U\times\mathbb{C}^{n}$.
+It follows that the set $F(K)$ is holomorphically convex in $U\times\mathbb{C}^{n}$, and hence in $X\times\mathbb{C}^{n}$ since $U$ is Runge in $X$.
+Let $f$ be a holomorphic function on a Stein neighbourhood $V\subset\Omega$ of $K$.
+We may assume that $\pi(V)\subset U$.
+Then, $\tilde{f}=f\circ F^{-1}:F(V)\to\mathbb{C}$ is a holomorphic function on the open subset $F(V)$ in the closed complex submanifold $F(\Omega_{U})$ of $U\times\mathbb{C}^{n}$.
+By the Cartan–Oka extension theorem, $\tilde{f}$ extends to a holomorphic function on a neighbourhood of $F(K)$ in $U\times\mathbb{C}^{n}$.
+(This can also be seen by precomposing $\tilde{f}$ by a holomorphic retraction from a neighbourhood of $F(\Omega_{U})$ in $U\times\mathbb{C}^{n}$ onto $F(\Omega_{U})$, given by Docquier and Grauert [14]; see also [19, Theorem 3.3.3, p.
+74].) By the Oka–Weil theorem, $\tilde{f}$ can be approximated uniformly on $F(K)$ by functions $g\in\mathscr{O}(X\times\mathbb{C}^{n})$.
+Every such function is of the form $g(x,w)=\sum_{\alpha\in\mathbb{Z}_{+}^{n}}g_{\alpha}(x)w^{\alpha},$ where $w=(w_{1},\ldots,w_{n})$ are coordinates on $\mathbb{C}^{n}$ and $g_{\alpha}\in\mathscr{O}(X)$ for every $\alpha\in\mathbb{Z}_{+}^{n}$.
+By cutting off the power series at a finite level gives functions $\mathscr{A}(X\times\mathbb{C}^{n})$ (polynomials in $w$ depending holomorphically on $x\in X$) approximating $g$ on $F(K)$.
+Precomposing such functions with $F$ gives the desired approximation of $f$ by functions in $\mathscr{A}(\Omega)$.
+∎
+The following is an immediate corollary to Theorem 3.4.
+Corollary 3.5.
+(Assumptions as in Theorem 3.4.) A compact set $K$ in $\Omega$ is $\mathscr{O}(\Omega)$-convex if and only if it is $\mathscr{A}(\Omega)$-convex.
+It follows that the functions in $\mathscr{A}(\Omega)$ satisfy the axioms of a Stein manifold, i.e., they separate points in $\Omega$ and provide holomorphic convexity.
+4. A relative Oka principle for fibrewise algebraic maps
+In this section, we describe a relative Oka principle (see Theorem 4.1) for holomorphic fibrewise algebraic maps from the Teichmüller family $V(g,n)$ $(n\geq 1)$, or more generally from any manifold $\Omega=Z\setminus H$ as in Theorem 3.1, to a class of complex algebraic manifolds which includes all flexible manifolds in the sense of Arzhantsev et al.
+[9].
+In order to state this result, we must recall the following notions; see Gromov [29, 0.5, p.
+8.5.5], [19, Definition 5.6.13], or [20, Definition 3.1].
+A dominating holomorphic spray on a complex manifold $Y$ is a holomorphic map $s:E\to Y$ from the total space of a holomorphic vector bundle $E\to Y$ such that
+(4.1) $s(0_{y})=y\ \ \text{and}\ \ ds_{0_{y}}(E_{y})=T_{y}Y\ \ \text{for every $y\in Y$.}$
+Here, $0_{y}$ is the origin of the fibre $E_{y}$ of $E$ over the point $y\in Y$, and $E_{y}$ is considered as a $\mathbb{C}$-linear subspace of the tangent space $T_{0_{y}}E$.
+(Identifying $Y$ with the zero section of $E$, we have $TE|Y\cong TY\oplus E$.) The spray is algebraic if the vector bundle $E\to Y$ and the map $s:E\to Y$ are algebraic.
+A complex manifold $Y$ is elliptic if it admits a dominating holomorphic spray, and is special elliptic if such a spray exists on a trivial bundle $E=Y\times\mathbb{C}^{m}$ for some $m\in\mathbb{N}$.
+An algebraic manifold $Y$ is algebraically (special) elliptic if these conditions hold with algebraic sprays.
+An algebraic manifold $Y$ is said to be flexible [9] if its tangent bundle $TY$ is pointwise spanned by finitely many LND’s, that is, algebraic vector fields $V_{j}$ with complete algebraic flows $\phi_{t}^{j}$ $(t\in\mathbb{C},\ j=1,\ldots,m)$.
+The map $s:Y\times\mathbb{C}^{m}\to Y$ defined by
+$s(y,t_{1},\ldots,t_{m})=\phi^{1}_{t_{1}}\circ\cdots\phi^{m}_{t_{m}}(y)$
+is a dominating algebraic spray on $Y$, so a flexible manifold is algebraically special elliptic. References to examples of flexible manifolds can be found in [20, p. 394].
+The following result is an analogue of the relative Oka principle for algebraic maps from affine algebraic varieties to algebraically elliptic manifolds; see [18, Theorem 3.1], [19, Theorem 6.15.1], and [20, Theorem 6.4].
+Theorem 4.1.
+Let $\pi:Z\to X$, $H\subset Z$, and $\Omega=Z\setminus H$ be as in Theorem 3.1, and let $Y$ be an algebraically special elliptic manifold.
+Given a map $f:Z\to Y$ of class $\mathscr{A}(Z,Y)$ (i.e., $f$ is holomorphic and fibrewise algebraic), a compact $\mathscr{O}(\Omega)$-convex set $K\subset\Omega$, and a homotopy of holomorphic maps $f_{t}:U\to Y$ $(t\in[0,1)$ on an open neighbourhood $U$ of $K$ with $f_{0}=f|_{U}$, there are holomorphic maps $F:\Omega\times\mathbb{C}\to Y$ such that $F(\cdotp,0)=f$ and $F(\cdot,t)\in\mathscr{A}(\Omega,Y)$ approximates $f_{t}$ as closely as desired uniformly on $K$ and uniformly in $t\in[0,1]$.
+In particular, a holomorphic map $\Omega\to Y$ that is homotopic to a map in $\mathscr{A}(\Omega,Y)$ is a limit of maps in $\mathscr{A}(\Omega,Y)$ uniformly on compacts.
+This holds in particular for the Teichmüller family $\Omega=V(g,n)$ for any $g\geq 0$ and $n\geq 1$.
+Proof.
+It suffice to inspect the proof of [18, Theorem 3.1] and apply Theorem 3.4. Let $p:E\to Y$ be an algebraic vector bundle and $s:E\to Y$ be a dominating algebraic spray. Consider the commuting diagram
+[…]
+where the map $\tilde{p}$ in the first column is the $f$-pullback of the vector bundle $p:E\to Y$, the map $\iota$ is the top row is the natural inclusion, and $\tilde{s}=s\circ\iota$ is a dominating holomorphic spray over $f$ (the pullback of $s$ by $f$).
+Note that $\tilde{s}=f$ holds on the zero section of $f^{*}E$.
+Since $f$ is fibrewise algebraic and $s$ is algebraic, the maps $\iota$ and $\tilde{s}$ are algebraic on $f^{*}E|\Omega_{x}$ for every $x\in X$.
+Replacing the spray bundle $(E,p,s)$ with a sufficiently high iterate of itself (see [19, Definition 6.3.5] for this notion) and shrinking $U$ around $K$ if necessary, [19, Proposition 6.5.1] give a lift of the homotopy of holomorphic maps $f_{t}:U\to Y$ in the theorem to a homotopy of holomorphic sections $\xi_{t}:U\to f^{*}E|_{U}$, with $\xi_{0}$ the zero section, such that $\tilde{s}\circ\xi_{t}=f_{t}$ for $t\in[0,1]$.
+Assume now that $p:E\to Y$ is a trivial algebraic vector bundle.
+Then, every iterate of $(E,p,s)$ is also a trivial bundle.
+Since $f:\Omega\to Y$ is fibrewise algebraic, the pullback bundle $f^{*}E\to\Omega$ is fibrewise algebraically trivial.
+Hence, sections of $f^{*}E\to\Omega$ can be identified with maps from $\Omega$ to the fibre, which is a Euclidean space.
+Thus, Theorem 3.4 lets us approximate every section $\xi_{t}$ uniformly on $K$ by a holomorphic fibrewise algebraic section $\tilde{\xi}_{t}:\Omega\to f^{*}E$.
+The map $F_{t}=\tilde{s}\circ\tilde{\xi}_{t}:\Omega\to Y$ is then of class $\mathscr{A}(\Omega,Y)$ and it approximates $f_{t}$ on $K$ for every $t\in[0,1]$.
+The approximation can be made uniform in $t\in[0,1]$ by including $[0,1]$ in a standard way into $\mathbb{C}$ and applying Theorem 3.4 with the compact holomorphically convex set $K\times[0,1]$ in $\Omega\times\mathbb{C}$.
+(The homotopy $\{\xi_{t}\}_{t\in[0,1]}$ must first be approximated on $K\times[0,1]$ by a holomorphic map in a neighbourhood of $K\times[0,1]$ in $\Omega\times\mathbb{C}$, which is an application of Mergelyan’s theorem.) This gives a holomorphic map $F:\Omega\times\mathbb{C}\to Y$ satisfying the conclusion of the theorem which is algebraic on $\Omega_{x}\times\mathbb{C}$ for every $x\in X$.
+∎
+Remark 4.2.
+The relative algebraic Oka principle in [18, Theorem 3.1] holds for all algebraically elliptic target manifolds $Y$.
+(It was stated for the ostensibly bigger class of algebraically subelliptic manifolds, but it has recently been shown by Kaliman and Zaidenberg [36] that these two classes of manifolds coincide.
+See also [20, Theorem 6.2] for a more complete list of equivalent algebraic ellipticity conditions.) We do not know whether Theorem 4.1 holds for this bigger class of manifolds.
+An example of a flexible manifold, which is of particular interest in the theory of minimal surfaces in Euclidean spaces $\mathbb{R}^{n}$, $n\geq 3$, is the punctured null quadric
+$\mathbf{A}=\{z=(z_{1},\ldots,z_{n})\in\mathbb{C}^{n}\setminus\{0\}:z_{1}^{2}+z_{2}^{2}+\cdots+z_{n}^{2}=0\}.$
+(See [4, Proposition 1.15.3] and [8].) Algebraic 1-forms with values in $\mathbb{A}$ on finitely punctured compact Riemann surfaces, which satisfy the real period vanishing conditions on closed curves in $M$, gives rise to conformal minimal surfaces $M\to\mathbb{R}^{n}$ of finite total curvature via the Enneper–Weierstrass representation formula, and vice versa.
+(See [40] and [4, Chap.
+4] for a discussion of this topic.) Hence, the fact that the Teichmüller family $V(g,n)$ for $n\geq 1$ admits nontrivial fibrewise algebraic maps $V(g,n)\to\mathbf{A}$ raises the question whether there one can realise the fibres of $V(g,n)$ as a family of minimal surfaces of finite total curvature depending holomorphically on the parameter $t\in T(g,n)$.
 Acknowledgements.
 Research was supported by the European Union (ERC Advanced grant HPDR, 101053085) and grants P1-0291 and N1-0237 from ARIS, Republic of Slovenia.
-I wish to thank Finnur Lárusson for posing the question answered by Theorem 1 in a private communication.
+I wish to thank Finnur Lárusson for asking the question answered by Theorem 1.1 and for proposing Corollary 1.3 in a private communication, and Daniel Huybrechts and Vladimir Marković for useful remarks and suggestions.